# Tint Your Models Task-wise for Improved Multi-task Model Merging

### v1

@@ Abstract
Traditional model merging methods for multi-task learning (MTL) address task conflicts with straightforward strategies such as weight averaging, sign consensus, or minimal test-time adjustments. This presumably counts on the assumption that a merged encoder still retains abundant task knowledge from individual encoders, implying that its shared representation is sufficiently general across tasks. However, our insight is that adding just a single trainable task-specific layer further can bring striking performance gains, as demonstrated by our pilot study. Motivated by this finding, we propose Model Tinting , a new test-time approach that introduces a single task-specific layer for each task as trainable adjustments. Our method jointly trains merging coefficients and task-specific layers, which effectively reduces task conflicts with minimal additional costs. Additionally, we propose a sampling method that utilizes the difference in confidence levels of both merged and individual encoders. Extensive experiments demonstrate our method’s effectiveness, which achieves state-of-the-art performance across both computer vision and natural language processing tasks and significantly surpasses prior works. Our code is available at https://github.com/AIM-SKKU/ModelTinting.
@@ 1 Introduction
Multi-task learning (MTL) [67] aims to develop models capable of handling multiple tasks simultaneously, enhancing efficiency and enabling knowledge transfer across tasks. Recent advances have facilitated the merging of independently fine-tuned models into a unified framework, significantly improving performance. Inspired by weight-space averaging [59, 58], Task Arithmetic [19], which formalizes task-specific knowledge as task vectors — the difference between pre-trained and fine-tuned models — opened new horizons in MTL.
However, we argue that negative transfer between tasks, which is a challenge in traditional MTL, still persists in model merging. To mitigate task conflicts arising from model merging, recent methods [63, 10, 66, 18] have employed predefined strategies focused on reducing parameter interference in shared representations. While effective, these methods involve ad-hoc procedures and focus entirely on shared representations to capture task-specific characteristics. Relying solely on such fixed strategies could limit the preservation of task-specific knowledge, creating a bottleneck in model merging.
This raises two fundamental questions: 1) Is relying solely on predefined strategies using shared representations sufficient to preserve task-specific knowledge? 2) Is there an effective way to capture task-specific features without relying on complicated training methods? To answer these questions, we start with a pilot study revealing that pairing a single respective task-specific layer with different task encoders trained on different tasks consistently enhances cross-task performance. This suggests that adding a minimal trainable layer (e.g., the classifier head) that captures task-specific features complements the task-agnostic knowledge in shared representations. This finding became the foundation of our proposed method.
In this paper, we propose a novel test-time MTL merging method dubbed Model Tinting that introduces respective task-specific layers through minimal learnable adjustments. By refining each layer to focus on its respective knowledge, our approach captures distinct task details, allowing the merged model to extract relevant features while retaining shared representations. Unlike previous methods [64, 50, 65], the proposed method focuses on enhancing task-specific knowledge at the output level. Our approach jointly trains merging coefficients and task-specific layers without requiring explicit costly optimizations, as a single coefficient set and layer are assigned to each task. To prevent the merged model from adopting incorrect predictions from individual models, we further propose Model Tinting++ . This method introduces a sampling strategy that excludes instances where individual models are incorrect, but the merged model predicts correctly based on confidence.
We demonstrate Model Tinting through extensive experiments, which achieve state-of-the-art performance across both vision and NLP domains (partially visualized in Figure 1) for different model scales and task configurations. We also provide some justifications for our design and some insights drawn from the empirical studies. Beyond the performance gains, our work demonstrates that effective multi-task learning does not require heavy fine-tuning that disrupts shared representations. Instead, task-specific knowledge can be naturally preserved and enhanced through minimal task-specific adaptations.
@@ 2.2 Model Merging for Multi-task Learning
A foundational work in model merging for MTL is Task Arithmetic [19], which defines task vectors as the difference between fine-tuned and pre-trained models, enabling the combination of weights across multiple tasks through arithmetic operations. However, this method faces two main issues: parameter interference and maintaining task-specific information to prevent a merged model from becoming overly generalized. Recent approaches tackle interference through parameter drop, redundancy reduction, and test-time adjustments. For instance, Ties-Merging [63] addresses redundant parameters and sign conflicts; DARE [66] randomly removes redundancies and rescales key parameters; EMR-Merging [18] uses masks and rescaling to retain critical values; AdaMerging [65] learns the importance of task vectors at test time.
To preserve more task-specific details, WEMoE [50] and Twin-Merging [32] utilize trained routers to adjust shared and task-specific components, while Representation Surgery [64] refines the representation space with test-time training on unlabeled data. Overall, existing model merging methods assume shared features alone can capture all task-specific characteristics, but this often fails, losing critical task information. Our proposed method addresses this by focusing on task-specific layers rather than just shared representations. By refining these layers while keeping the encoder shared, Model Tinting captures essential task-specific features often overlooked in shared layers. This approach optimizes merging coefficients and task-specific layers, minimizing conflicts and often surpassing individual model performance.
3 Preliminary
3.1 Background
Problem definition. Multi-task Learning (MTL) [67] aims to train a single model to achieve strong performance across multiple tasks. Given the input $x^{k}\in\mathbb{R}^{d_{in}}$ for the kth task, a $L$-layer encoder $f(x^{k};\phi)$ with the weights $\phi=\{\phi^{1},\ldots,\phi^{L}\}$ outputs the representation $\hat{z}^{k}\in\mathbb{R}^{d_{out}}$. For each of the $K$ tasks, $\hat{z}^{k}$ is fed into a task-specific layer $g(\cdot\ ;\theta_{k})$, where $\theta_{k}$ represents weights unique to task $k$, to perform classification (e.g., the classification head in CLIP [43]). Each task output should align with the task nature, and an improved MTL model should provide higher accuracies across tasks. We employ two types of weights to address MTL: $\phi_{\text{pre}}$ and $\phi_{k}$, which denote a pre-trained weight and task-specific fine-tuned weights for task $k$, respectively. These weights form the basis of novel weight merging methods [58, 22] aimed at excelling in MTL.
@@ 3.1 Background
AdaMerging [65] improved the Task Arithmetic by building a merged model $f(x;\phi_{\text{MTL}})$ through learning $\lambda$ in a task-wise and layer-wise manner, defined as $\phi_{\text{MTL}}^{l}=\phi_{\text{pre}}^{l}+\sum_{k=1}^{K}\lambda_{k}^{l}\cdot\tau_{k}^{l}$. Each merging coefficient $\{\lambda_{k}^{1},\ldots,\lambda_{k}^{L}\}$ is trained by optimizing the objective: $\underset{\{\lambda_{k}^{l}\}}{\min}\sum_{k,x^{k}\in\mathcal{X}}\mathcal{L}_{k}\left(g_{k}\left(f(x^{k};\phi_{\text{MTL}});\theta_{k}\right)\right).$ Without ground-truth labels (for a test set $\mathcal{X}$), AdaMerging adopts an entropy minimization objective $\mathcal{L}_{k}$ to learn each coefficient. Although AdaMerging effectively merges multi-task weights, we believe there is still room for improvement.
@@ 3.2 Pilot Study
The success of prior MTL merging methods [32, 66, 63, 65, 18] lies in reducing conflicts in the encoder’s representations across tasks. However, we argue that adopting the shared representations without task-specific knowledge may harm MTL performance due to the task-agnostic nature of the merged encoder [^1] Those MTL merging methods generally achieve overall improvements across tasks through weight averaging, which suggests that the merged encoder becomes more likely to be task-agnostic..
Study outline. Our goal is to validate the claim and take it a step further by showing that the following solution works: a trained task-specific layer in a merged model could enhance task-specific capabilities in task-agnostic representation. Here, we use the classifier as the layer for analysis, as it is the most straightforward option. If our claim is correct, the merging methods [65, 64] could benefit performance-wise from minimal additional training.
Cross-task evaluation. Given that each $k$-th task has its own fine-tuned model, $g_{k}^{bs}\left(f(x^{k};\phi_{k});\theta_{k}^{bs}\right)$, which serves as our control group, where $g_{k}^{bs}(\cdot;\theta_{k}^{bs})$ denotes the base (or primitive) classifier in the pre-trained model (e.g., CLIP [43]). We define the cross-task performance, computed by using the visual encoder from another task $m$ in the form $g_{k}^{bs}\left(f(x^{k};\phi_{m});\theta_{k}^{bs}\right)$ for a given task $k$ (i.e., swapping the trained layer only with the original one). We emulate the optimization-based merging [65] to observe whether the trained weight would steer towards a task-specific direction. To this end, a weight-averaged encoder [19], $f(x^{k};\phi_{\text{MTL}})$ is first employed as the backbone to generate task-agnostic representations. Training is performed to minimize the following objective with respect to the designated layers ($g_{k}^{tr}$)’s weight ($\theta_{k}^{tr}$):
@@
where $y^{k}$ denotes the corresponding ground-truth label for the input sample $x^{k}$. This setup uses labeled samples to confirm the encoder’s maximal capability by utilizing ground-truth labels. We then compute cross-task scores by pairing each trained classifier $g_{k}^{tr}(\cdot;\theta_{k}^{tr})$ with individual encoders, which is formulated as $g_{k}^{tr}(f_{m}(x^{k};\phi_{m});\theta_{k}^{tr})$ for all $m\neq k$.
(a) Primitive classifiers
(b) Trained classifiers
Figure 2: Cross-task evaluations across diverse encoders and heads from different tasks. Each cell $(i,j)$ displays the accuracy when features from task $i$’s visual encoder are classified by task $j$’s classifier. The upper value in a cell shows the absolute task accuracy, and the number below (in parentheses) denotes the relative performance compared to the diagonal. (a) Primitive (untrained) classifiers generally deteriorate when evaluated on other tasks or using different encoders. (b) Our trained classifiers show mostly reduced performance loss (i.e., higher relative numbers). This suggests that task-specific adjustments effectively refine the task-agnostic representations task-wise.
Our interpretation. Figure 2 shows cross-task evaluations via individual visual encoders with their respective primitive classifiers; Figure 2 shows results by combining the individual visual encoders with the trained classifiers. We observe consistent improvements in cross-task performance when using the classifiers trained on other tasks. For instance, integrating the MNIST classifier with the Cars encoder achieves 58.1% accuracy (Figure 2), surpassing the base classifier’s 51.8% (Figure 2). These results suggest that training the classifier on merged features for the MNIST task also facilitates the alignment of individual Cars features with the MNIST task. This demonstrates the trained layer’s enhanced capability to refine task-specific outputs from task-agnostic representations. We further conjecture that an intermediate layer in the backbone could function similarly, and this will also be empirically proven by our experiments (see Figure 4).
A remaining challenge is that weight merging is typically unavailable to access the ground-truth labels for queries in practice. Therefore, our method is designed with this constraint in mind so that it can work in real-world scenarios.
@@ 4 Method
This section introduces our method, inspired by the pilot study, which incorporates task-specific adaptation during weight merging to improve MTL performance significantly.
@@ 4.1 Model Tinting
Motivation. Our pilot study revealed that training a single layer (e.g., a classifier head) on merged representations enhances task-specific capabilities, potentially boosting overall MTL performance. Motivated by this finding, we explore an effective merging method that selectively adapts a few layers in the merged encoder to each task during the merging process. We conjecture that this would merit improved MTL performance by injecting task-specific knowledge into the task-agnostic encoder, barely incurring extra training costs.
Leveraging task-specific knowledge at test-time. We design our method based on test-time approaches [65, 64], where ground-truth labels are unavailable. A straightforward approach would be to adopt the entropy [65] as the metric replacing any supervisory signals. Since the entropy roughly follows the cross-entropy’s trend [65, 38], we believe that the self-entropy may not fully align when data cardinality is insufficient. We empirically found that the entropy does not closely approximate the cross-entropy (with ground-truths) during training (see Table 1), so we identify an effective way of employing the cross-entropy.
Drawing from what we found in the pilot study, we mimic each task encoder’s knowledge to equip the classifier with task-specific knowledge. We define an objective to align the distribution of each fine-tuned individual model, $C_{k}^{\text{ft}}=\{g_{k}^{\text{bs}}\left(f(x^{k};\phi_{k});\theta_{k}^{\text{bs}}\right)|{x^{k}{\in}\mathcal{X}}\}$, with the merged model’s, $C_{k}^{\text{merged}}=\{g_{k}^{\text{tr}}\left(f(x^{k};\phi_{\text{MTL}});\theta_{k}^{\text{tr}}\right)|{x^{k}{\in}\mathcal{X}}\}$ on test data $\mathcal{X}$. While some semi-supervised techniques [68, 51, 41, 1] could be employed, we opt for a simpler approach [29]: training the merged model to match the fixed output distributions of the individual models [^2] While individual models could be updated during training or replaced with other merged models as alternative options, we opt for a simpler approach. We provide the related study in Figure 6. using the cross-entropy loss. This optimization objective is defined as follows:
@@
where $\mathcal{L}_{k}$ denotes the loss for each task, with $g_{k}^{\text{bs}}$ representing the primitive task-specific layers, and $g_{k}^{\text{tr}}$ representing the learnable task-specific layers. Notably, in contrast to AdaMerging [65], our approach jointly optimizes both the merging coefficients and the task-specific layers rather than focusing solely on the merging coefficients.

### v2

# Why Train Everything? Tint a Single Layer for Multi-task Model Merging

@@ Abstract
Model merging integrates independently fine-tuned models into a single multi-task model, offering a flexible alternative to joint training. However, many existing model merging methods introduce additional task-specific components, increasing complexity and requiring extra modifications. We propose Model Tinting , a lightweight yet highly effective approach that improves model merging by updating just a single layer, accounting for as low as 0.5% of total parameters. Our key observation is that explicit task-specific modules are not necessary; instead, subtle adjustments to a single layer can effectively capture task-specific variations within the merged model while maintaining generalization. We introduce a confidence-based filtering mechanism to alleviate the impact of unreliable predictions from individual models on the merged model. Extensive experiments across vision and NLP tasks demonstrate that Model Tinting achieves state-of-the-art performance, even in challenging dense prediction tasks. Our code is available at https://github.com/AIM-SKKU/ModelTinting.
@@ 1 Introduction
Multi-task learning (MTL) [4] enables models to handle multiple tasks efficiently by facilitating knowledge transfer. Model merging [67, 22, 68] emerged as an alternative to conventional MTL [76, 40, 28, 20] by combining independently fine-tuned models into a single model. This approach allows for flexible adaptation, reduces computational overhead, and better preserves task-specific knowledge. A key challenge in model merging is balancing generalization and task specialization. On one hand, a merged model should retain shared knowledge across tasks [72, 74], but on the other hand, it must preserve task-specific adaptations that are crucial for individual task performance [73, 21].
Early methods [23, 72, 75, 74, 71] primarily focused on constructing task-agnostic models by averaging weights or aligning parameters across models, assuming that similar models in weight space would naturally form a meaningful shared representation. However, this generalization-first strategy often led to a loss of task-specific information, reducing performance on individual tasks. This issue became more pronounced when merging models trained on diverse or conflicting tasks, as simple averaging often failed to preserve specialized adaptations essential for high accuracy.
Recent approaches [21, 73, 58, 36] explored task-specific parameterization, with some incorporating additional components, in some cases, are trainable to enhance adaptation. These methods typically added task-specific modules or learnable merging coefficients, selectively adjusting how different tasks influence the final model. While this enhanced task adaptation, it also introduced challenges such as increased model complexity, higher computational costs in both training and inference, and scalability issues when applied to many tasks. Furthermore, many merging strategies still relied on a fixed shared model [23, 72, 75, 11, 36, 21, 73, 58], often formed by merging encoders without adaptation. As a result, the merged model struggled to retain fine-grained task distinctions and failed to effectively leverage task-relevant cues, leading to suboptimal performance in diverse task distributions (see Figure 1). These challenges raise two fundamental questions: 1) Can task-specific features be captured without introducing additional modules? 2) Is a predefined task-agnostic backbone sufficient to retain task-specific knowledge effectively?
In this paper, we propose Model Tinting , a simple yet effective model merging method that preserves task-specific knowledge with minimal overhead. Instead of introducing additional modules, our method maintains task-specific information through minimal trainable adjustments, updating a single layer that accounts for fewer than 0.5% of parameters in ViT-B/32. By jointly optimizing task-specific layers and merging coefficients, it refines both components under the guidance of individually fine-tuned models. To further enhance reliability, we introduce a confidence-based filtering mechanism that prevents the model from inheriting incorrect predictions from individual experts. By selectively removing unreliable supervision signals, this mechanism improves robustness and enables the merged model to sometimes surpass individual task models.
Pub. Method Task-specific Adaptation No Extra Module Adaptive Shared Model
ICLR 2023 Task Arithmetic [23]
ICLR 2024 AdaMerging [74]
ICML 2024 Surgery [73]
ICML 2024 WEMoE [58]
NeurIPS 2024 EMR-Merging [21]
NeurIPS 2024 TWIN-Merging [36]
Model Tinting
Table 1: Comparison of recent model merging methods.
Table 1 shows the main differences between existing methods and our method. Model Tinting requires no additional modules beyond a single trainable layer per task, making it lightweight and easy to deploy. Furthermore, we also adjust the task-agnostic model (i.e., shared model) to better leverage the effect of the task-specific layer, further enhancing adaptation for each task. Through extensive experiments across image classification, dense prediction, and NLP tasks, we demonstrate that Model Tinting achieves state-of-the-art performance across different model scales and task configurations. Notably, our method excels in dense prediction tasks, a challenging yet underexplored domain in prior model merging studies. Our approach achieves remarkably strong results, demonstrating its ability to handle complex task structures effectively. Our main contributions are as follows:
  • We propose a lightweight MTL model merging method that preserves task-specific knowledge with minimal trainable adjustments, requiring no additional modules.
  • To improve robustness, we introduce a confidence-based filtering mechanism that selectively removes unreliable supervision signals from individual models, preventing error propagation in the merged model.
  • Our method achieves state-of-the-art performance across diverse tasks, including image classification, dense prediction, and NLP, particularly excelling in dense prediction, an underexplored area in model merging.
@@ 2.2 Model Merging for Multi-task Learning
Negative transfer remains a fundamental challenge in model merging for MTL, as merging often inherits task interference within the shared representation. The goal is to construct a single model that generalizes well across all tasks while minimizing such interference. Existing weight alignment methods [71, 55, 1] enable merging fine-tuned weights obtained from different pre-trained models. However, these approaches face scalability limitations as the number of tasks increases. A foundational work in model merging for MTL is Task Arithmetic [23], which introduces task vectors, defined as the difference between pre-trained and fine-tuned models. This enables arithmetic-based weight combinations across multiple tasks, allowing the merging of a large number of individual models. Subsequent methods refine this approach by explicitly addressing task conflict within the merged model. Ties-Merging [72] mitigates redundancy and sign conflicts, while DARE [75] removes redundant components and rescales key parameters. AdaMerging [74] learns task vector importance at test time, and LiNeS [64] reduces the impact of shallow layers to preserve general representations.
While merged models have traditionally been considered task-agnostic, recent findings [63, 19] suggest that they inherently encode task-specific knowledge. Motivated by these observations, recent approaches incorporate task-specific modules into the merged encoder to better align with the performance of individual models. Surgery [73] refines merged encoders by introducing task-specific adapters, while EMR-Merging [21] applies task-specific masking to a pre-defined merged model. Router-based methods [58, 36] determine task vector coefficients via routing mechanisms that depend on each task’s input. However, these methods introduce additional parameters and often overlook the role of a well-structured shared expert in maximizing task-specific module effectiveness. Model Tinting explicitly refines both shared and task-specific layers in a complementary manner, which allows it to easily capture hidden task-specific knowledge. This balanced design improves task adaptation and often surpasses individual model performance.
3 Background
3.1 Preliminary
Problem definition. Multi-task Learning (MTL) [4] aims to train a single model to achieve strong performance across multiple tasks. Given the input $x^{k}\in\mathbb{R}^{d_{in}}$ for the kth task, a $L$-layer encoder $f(x^{k};\phi)$ with the weights $\phi=\{\phi^{1},\ldots,\phi^{L}\}$ outputs the representation $\hat{z}^{k}\in\mathbb{R}^{d_{out}}$. For each of the $K$ tasks, $\hat{z}^{k}$ is fed into a task-specific layer $g(\cdot\ ;\theta_{k})$, where $\theta_{k}$ represents weights unique to task $k$, to perform classification (e.g., the classification head in CLIP [48]). We employ two types of weights to address MTL: $\phi_{\text{pre}}$ and $\phi_{k}$, which denote a pre-trained weight and task-specific fine-tuned weights for task $k$, respectively. These weights form the basis of novel weight merging methods [67, 26] aimed at excelling in MTL.
@@ 3.1 Preliminary
AdaMerging [74] improves Task Arithmetic by constructing a merged model $f(x;\phi_{\text{MTL}})$ through learning $\lambda$ in a task- and layer-wise manner, defined as $\phi_{\text{MTL}}^{l}=\phi_{\text{pre}}^{l}+\sum_{k=1}^{K}\lambda_{k}^{l}\cdot\tau_{k}^{l}$. Each merging coefficient $\{\lambda_{k}^{1},\ldots,\lambda_{k}^{L}\}$ is trained by optimizing: $\underset{\{\lambda_{k}^{l}\}}{\min}\sum_{k,x^{k}\in\mathcal{X}}\mathcal{L}_{k}\left(g_{k}\left(f(x^{k};\phi_{\text{MTL}});\theta_{k}\right)\right).$ Without ground-truth labels (for test set $\mathcal{X}$), AdaMerging adopts an entropy minimization objective $\mathcal{L}_{k}$ to learn each coefficient.
Figure 2: Target and cross-task evaluations across 8 tasks using ViT-B/32 with base and trained classifiers. We also evaluate both Task Arithmetic and AdaMerging encoders. Training classifiers on merged representations improves cross-task generalization while preserving target task performance. Additionally, performance varies depending on the shared encoder’s ability to retain task-specific features.
@@ 3.2 Pilot Study
The success of recent MTL merging methods [36, 21, 58, 73] lies in introducing additional task-specific components to extract hidden task-specific knowledge [19, 63] from a predefined merged encoder (i.e., shared encoder) [^1] Those MTL merging methods generally achieve overall improvements across tasks through weight averaging, which suggests that the shared encoder becomes more likely to be task-agnostic.. Instead, we argue that selecting appropriate layers within the model is sufficient without introducing additional modules. Furthermore, we suggest that it is crucial to adjust the shared encoder to better capture task-specific information.
Study outline. Our goal is to validate this claim and extend it further by demonstrating that a trained task-specific layer can enhance task-specific capabilities even within a task-agnostic representation. Additionally, we investigate whether the shared encoder affects the extraction of task-specific knowledge. We use the classifier as our task-specific layer for analysis due to its simplicity. If our hypothesis holds, merging methods [74, 73] could achieve improved performance with minimal additional training.
Task-specific training. We generate task-agnostic representations from the fixed merged encoder $f(x^{k};\phi_{\text{MTL}})$, which is obtained using Task Arithmetic [23] or AdaMerging [74]. Then, we train a base classifier $g_{k}(\cdot;\theta_{k})$ from the pre-trained model for the corresponding $k^{th}$ task (e.g., $k=1,...,8$). The following objective is minimized with respect to the weights ($\theta_{k}^{tr}$) of the selected layer ($g_{k}^{tr}$) during training:
@@
where $g_{k}^{tr}$ represents the trainable classifier and $y^{k}$ denotes the corresponding ground-truth label for the input sample $x^{k}$. This setup uses labeled samples to confirm the encoder’s maximal capability by utilizing ground-truth labels.
Target & cross-task evaluation. We evaluate each base $g_{k}^{bs}$ and trained classifier $g_{k}^{tr}$ using features from all individual encoders $f(\cdot;\phi_{\text{m}})$ (e.g., $648^{2}=$ 64 cases). To assess its effectiveness, we compute the average performance in two settings: 1) target task evaluation, where the classifier is tested on features from its corresponding encoder (i.e., $m=k$), and 2) cross-task evaluation, where it is tested on features from individual encoders trained on distinct tasks (i.e., $m\neq{}k$).
Interpretation. Figure 2 shows the average performance of target task and cross-task evaluations separately, using base classifiers and trained classifiers. We observe consistent improvements in cross-task performance when using the classifiers trained on other tasks. Detailed performance for each task pair is provided in Appendix C.1. These results suggest that training the classifier on merged features also facilitates the alignment of individual features with other tasks. This demonstrates the trained layer’s enhanced capability to refine task-specific outputs from task-agnostic representations. We further argue that an intermediate layer in the backbone could function similarly, and this will also be empirically proven by our experiments (see Figure 4).
Additionally, the choice of the shared encoder used for classifier training also affects performance. Interestingly, we observe that when using the Task Arithmetic features, cross-task performance improves significantly more than with the AdaMerging features. This highlights the importance of identifying a suitable shared encoder that better leverages task-specific knowledge. A remaining challenge is that weight merging is typically unavailable to access the ground-truth labels in practice. Therefore, our method is designed to handle this constraint and remains effective in real-world scenarios.
@@ 4.1 Model Tinting
Motivation. Our pilot study revealed that training a single layer (e.g., a classifier head) on merged representations enhances task-specific performance and can improve overall MTL effectiveness. Building on this finding, we propose an efficient merging method that fine-tunes a selected single layer in the merged encoder for each task, extending beyond standard adaptation approaches, while also refining the merged model.
Task-specific adaptation at test time. Our method follows test-time adaptation strategies [74, 73], where ground-truth labels are unavailable. A common approach is to use entropy [74] as a proxy for supervision. However, entropy only roughly follows the trend of cross-entropy [74, 44] and may not align well when data cardinality is insufficient. Our empirical analysis (see Table 2) shows that entropy does not consistently approximate cross-entropy with ground-truth supervision during training. Thus, we propose a more effective way to incorporate cross-entropy.
Building on our pilot study (refer to Section 3.2), we align the classifier’s output with each task encoder’s knowledge to inject task-specific information. We define an objective to align each fine-tuned model’s distribution with that of the merged model on the test set $\mathcal{X}$. Formally, for each task $k$, we minimize the discrepancy between $C_{k}^{\text{ft}}=\{g_{k}^{\text{bs}}(f(x^{k};\phi_{k});\theta_{k}^{\text{bs}})\mid x^{k}\in\mathcal{X}\}$ and $C_{k}^{\text{merged}}=\{g_{k}^{\text{tr}}(f(x^{k};\phi_{\text{MTL}});\theta_{k}^{\text{tr}})\mid x^{k}\in\mathcal{X}\}$.
Instead of relying on semi-supervised techniques [77, 59, 46, 2], we adopt a simpler approach [33]: training the merged model to match the fixed output distributions of the individual models [^2] While individual models could be updated during training or replaced with other merged models as alternative options, we opt for a simpler approach. We provide the related study in Figure 5. using the cross-entropy loss. This optimization objective is defined as follows:
@@
where $\mathcal{L}_{k}$ represents the loss for each task, $g_{k}^{\text{bs}}$ are the base task-specific layers, and $g_{k}^{\text{tr}}$ are the learnable task-specific layers. Unlike AdaMerging [74], which optimizes only merging coefficients in the shared encoder, our approach jointly optimizes both coefficients and task-specific layers to better tailor representations for distinct tasks.
@@
Loss correlation. To validate our approach, we analyze how well different loss functions correlate with ground-truth-based cross-entropy loss, which serves as a strong supervision signal when labels are available. We compute the Spearman correlation [^3] Spearman correlation quantifies monotonic relationships and is often used for ranked data or non-linear associations. between model-generated losses and the ground-truth-based cross-entropy loss. A higher correlation indicates that the loss function provides a reliable training signal, even in the absence of explicit labels.
Table 2 shows that entropy initially correlates moderately with supervised cross-entropy. However, after training, this correlation significantly weakens, and in some cases, entropy diverges sharply (e.g. Cars: -1.19 drop). This suggests that entropy is not a stable proxy for supervision and may fail to maintain consistency as training progresses. This finding aligns with prior studies [74, 44], which observed that entropy-based losses can misrepresent model confidence, leading to unreliable optimization. In contrast, our model-generated cross-entropy loss starts with a higher correlation and maintains strong alignment with supervised cross-entropy throughout training. Unlike entropy, which fluctuates unpredictably, our approach preserves a stable and reliable training signal across tasks.
4.2 Model Tinting++
Refining Model Tinting . Model-generated supervisory signals contribute to performance improvement, but not all signals are equally beneficial. In particular, cases where the merged model correctly classifies a sample while individual models fail may introduce unreliable training data. To address this, we assume that only high-confidence samples from individual models should be used for training. Prior studies [16] suggest that well-calibrated models mitigate overconfidence and achieve greater robustness. Following this principle, confidence-aware filtering is expected to enhance the stability of the merged model [44].
Directly setting a fixed confidence threshold is challenging and may introduce unwanted biases. Instead, we propose a relative confidence-based filtering strategy: We compare the top-1 probabilities of individual and merged models and exclude samples where the merged model assigns a higher probability than the individual models. This ensures that the merged model receives supervision primarily from reliable samples, leading to more stable adaptation. As discussed in [43], filtering uncertain cases enhances model robustness, which is expected to improve MTL performance across various tasks.
(a) Prediction discrepancy
(b) Overall difference
Prediction discrepancy analysis. To evaluate the impact of our approach, we analyze discrepancies between merged model predictions and those of individual models, which often serve as a performance upper bound in MTL. Figure 3 presents two sets of bar graphs illustrating prediction inconsistencies. The upper bars represent samples correctly classified by individual models but misclassified by the merged model, while the lower hatched bars indicate cases where individual models fail but the merged model succeeds. A method with lower upper bars and higher lower bars is preferable, as it signifies reduced merging-induced errors and improved generalization.
Among the evaluated approaches, our method achieves the smallest prediction discrepancy, significantly reducing the number of misclassified samples. This suggests that the filtering strategy effectively retains the strengths of individual models while mitigating degradation caused by merging. Figure 3 further summarizes the overall impact by computing the net difference between the upper and lower bars. A smaller value in this metric indicates a more balanced model that preserves individual model accuracy while capturing additional gains. The proposed method consistently outperforms alternatives, demonstrating its effectiveness in minimizing errors while improving generalization, leading to enhanced MTL performance.